# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

### v1

@@ Abstract
With the increasing collection of time series data from various domains, there arises a strong demand for general time series forecasting models pre-trained on a large number of time-series datasets to support a variety of downstream prediction tasks. Enabling general time series forecasting faces two challenges: how to obtain unified representations from multi-domian time series data, and how to capture domain-specific features from time series data across various domains for adaptive transfer in downstream tasks. To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting. ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic and periodic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains. We also equip ROSE with a Time Series Register, which learns to generate a register codebook to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks. After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 8 real-world benchmarks. Remarkably, even in few-shot scenarios, it demonstrates competitive or superior performance compared to existing methods trained with full data.
@@ 1 Introduction
Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics [1]. However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity [2]. One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios. However, applying this paradigm to heterogeneous time series data presents challenges.
@@
Obtaining a unified representation from time series data across various domains is challenging. Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other [3], which is frequency superposition. As shown in Figure 1(a), different frequency components contain distinct semantic information. For example, low and high-frequency components represent long-term trends and rapid variations, respectively [4]. Additionally, different frequency components also contain cross-period information, revealing the interactions between short and long-term patterns in the time series. Thus, the frequency superposition leads to coupled semantic and periodic information in the time domain. Furthermore, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity. Existing pre-training frameworks [5, 6, 7], such as mask modeling and contrastive learning, were proposed to learn a unified representation from time series data. However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
@@
Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge. Multi-source time series data originate from various domains [8], whose data exhibit domain-specific information [2]. Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task [9]. However, as shown in Figure 1(a), existing time series pre-training frameworks [8, 2, 10] focus mainly on capturing generalized features during pre-training and overlook domain-specific features. Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks, as shown in Figure 1(b). Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer. Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training. 2) adaptive use of domain-specific information in various downstream tasks.
@@
To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE). First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic and periodic information. We decompose individual time series using the Fourier transform with a novel frequency-based mask method, and then convert it back to the time domain to obtain decoupled time series for reconstruction. It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations. Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data. By setting up a register codebook, we generate register tokens to learn each domain-specific information during pre-training. In a downstream scenario, the model adaptively selects vectors from the register codebook that are close to the target domain of interest. During fine-tuning, we incorporate learnable vectors into the selected register tokens to complement target specific information to perform more flexible adaptive transfer. As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks. The contributions are summarized as follows:
  • We propose ROSE, a novel general time series forecasting model using multi-domain datasets for pre-training and improving downstream fine-tuning performance and efficiency.
  • We propose a novel Decomposed Frequency Learning that employs multi-frequency masking to decouple complex temporal patterns, empowering the model’s generalization capability.
  • We propose TS-Register to capture domain-specific information in pre-training and enable adaptive transfer of target-oriented specific information for downstream tasks in fine-tuning.
  • Our experiments with 8 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
@@ 2 Related work
2.1 Time series forecasting
Traditional time series prediction models like ARIMA [11], despite their theoretical support, are limited in modeling nonlinearity. With the rise of deep learning, many RNN-based models [12, 13, 14] have been proposed, modeling the sequential data with an autoregressive process. CNN-based models [15, 16] have also received widespread attention due to their ability to capture local features. MICN [17] utilizes TCN to capture both local and global features, while TimesNet [18] focuses on modeling 2D temporal variations. However, both RNNs and CNNs struggle to capture long-term dependencies. Transformer-based models [3, 6, 19, 20, 21], with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction. However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
2.2 Time series pre-training
Masked modeling [5, 22, 23] and contrastive learning [24, 25, 26] have attracted significant attention in the time series domain. PatchTST [6] proposes to capture local semantic information by predicting masked subsequence patches. LaST [27] explores using variational inference to decompose time series into representations of seasonality and periodicity within a latent space. PITS [7] combines masked modeling and contrastive learning to obtain better generalization representations. However, the above methods are limited to one-to-one pre-training and fine-tuning, overlooking the frequency diversity and complexity within multi-domain data, thus making them challenging to directly apply to pre-training for multi-domain time series.
Multi-source time series pre-training has recently garnered widespread attention [28, 29, 30]. MOMENT [31] and MOIRAI [8] adopt a BERT-style pre-training approach, while Timer [2] and PreDcT [32] use a GPT-style pre-training approach, giving rise to improved performance in time series prediction. However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models. Different from previous approaches, ROSE pre-trains on large-scale data from various domains and it considers both generalized representations and domain-specific information, which facilitates flexible adaptive transfer in downstream tasks.
@@ 3.1 Architecture
The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning. First, the model is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction. We set up the self-supervised reconstruction task to help the model understand time series comprehensively and set the supervised prediction task to enhance the model’s few-shot learning ability. Then, ROSE fine-tunes with a target dataset in the downstream scenario.
As shown in Figure 2, ROSE employs the Encoder-Decoder architecture to model time series. The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies [33]. ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting [6].
Input Representations. To enhance the generalization of ROSE for adaptive transferring from multi-domains to different target domains, we model the inputs $\mathbf{x}$ with patch tokens and register tokens. Patch tokens are obtained by partitioning the time series using patching layers [6], to preserve local temporal information. Register tokens are obtained by linear mapping and clustering each of the entire time series input into discrete embedding, to capture domain-specific information, which will be introduced in Section 3.3.
@@ 3.2 Decomposed frequency learning
As shown in Figure 1, time series data are composed of multiple combined frequency components, resulting in overlapping of different temporal variations. Furthermore, low-frequency components typically contain information about overall trends and longer-scale variations, and high-frequency components usually contain information about short-term fluctuations and shorter-scale variations, therefore, understanding time series from low and high frequencies separately benefits general time series representation learning. Based on the observations above, we propose decomposed frequency learning based on multi-freq masking to understand the time series from multiple-frequency perspectives, which enhances the model’s ability to learn generalized representations.
Figure 3: An illustration of decomposed frequency learning. Based on the pre-defined thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
Multi-freq masking. Given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) [34] to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{R}^{L/2}$.
@@
To separately model high-frequency and low-frequency information in time series, we sample $K_{\mathrm{f}}$ thresholds ${\tau_{1},\tau_{2},\tau_{3},...,\tau_{K_{\mathrm{f}}}}$ and $K_{\mathrm{f}}$ random numbers ${\mu_{1},\mu_{2},\mu_{3},...,\mu_{K_{\mathrm{f}}}}$ for multi-frequency masks, where $\tau\in\mathrm{Uniform}(0,a)$, $a<L/2$ , and $\mu\in\mathrm{Bernoulli}(1,p)$. Each pair of $\tau_{i}$ and $\mu_{i}$ corresponds to the $i_{th}$ frequency mask. This generates a mask matrix $\mathbf{M}\in\{0,1\}^{K_{\mathrm{f}}\times L/2}$, where each row corresponds to the $i_{th}$ frequency mask, each column corresponds to the $j_{th}$ frequency, and each element $m_{ij}$ is 0 or 1, meaning that the $j_{th}$ frequency is masked with the $i_{th}$ frequency mask or not.
$m_{ij}=\left\{\begin{aligned} &\quad\mu_{i}&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}<\tau_{i}\\ &(1-\mu_{i})&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}>\tau_{i}\end{aligned}\right.,$ (2)
where $\tau_{i}$ and $\mu_{i}$ denote the threshold and random number for the $i_{th}$ frequency domain mask, $\mathbf{x}_{\mathrm{freq}_{j}}$ represents the $j_{th}$ frequency of $\mathbf{x}_{\mathrm{freq}}$. If $\mu_{i}=1$, it indicates masking of the high frequencies, whereas if $\mu_{i}=0$, it signifies the masking of the low frequencies.
After obtaining the mask matrix $\mathbf{M}$, we replicate $\mathbf{x}_{\mathrm{freq}}$ $K_{\mathrm{f}}$ times to get the $\mathbf{X}_{\mathrm{freq}}\in\mathbb{R}^{K_{\mathrm{f}}\times L/2}$ and perform element-wise Hadamard product with the mask matrix $\mathbf{M}$ to get masked frequency of time series. Then, we use the inverse Real Fast Fourier Transform (irFFT) to convert the results from the frequency domain back to the time domain and get $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}=\{\mathbf{x}_{\mathrm{mask}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$, where each $\mathbf{x}_{\mathrm{mask}}^{i}\in\mathbb{R}^{L}$ corresponding to masking with a different threshold $\tau_{i}$.
@@
Representation learning. After obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token. We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3. Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the transformer encoder to obtain the representation of each masked series. These representations are then aggregated to yield a unified representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
$\mathbf{S}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
Reconstruction task. After obtaining the representation $\mathbf{S}$, we feed it into the reconstruction decoder and ultimately reconstruct the original sequence $\mathbf{\hat{x}}\in\mathbb{R}^{L}$ through the reconstruction head. As frequency domain masking affects the overall time series, we compute the Mean Squared Error (MSE) reconstruction loss for the entire time series.

### v2

# ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning

@@ Abstract
With the increasing collection of time series data from various domains, there arises a strong demand for general time series forecasting models pre-trained on a large number of time-series datasets to support a variety of downstream prediction tasks. Enabling general time series forecasting faces two challenges: how to obtain unified representations from multi-domian time series data, and how to capture domain-specific features from time series data across various domains for adaptive transfer in downstream tasks. To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting. ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains. We also equip ROSE with a Time Series Register, which learns to generate a register to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks. After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 7 real-world benchmarks. Remarkably, it demonstrates competitive or superior few-shot and zero-shot abilities.
@@ 1 Introduction
Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics (Qiu et al., 2024). However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity (Liu et al., 2024). One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios or direct predict without fine-tuning. Following this idea, foundation models for time series forecasting have recently raised growing attention, continually scaling up the pre-training datasets and model sizes to improve the generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024). However, an excessively large scale also causes increasing costs in training and inference, which may go against the original intention of a general model, especially in resource-constrained situations. In addition to scaling up, general time series forecasting models can also be designed from the perspectives of pre-training tasks and downstream transfer adaptation. From these two angles, we identify the following challenges.
@@
Obtaining a unified representation from time series data across various domains is challenging. Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022), which is frequency superposition. As shown in Figure 1(a), different frequency components contain distinct semantic information. For example, low and high-frequency components represent long-term trends and rapid variations, respectively (Zhang et al., 2022). Furthermore, different datasets exhibit diverse frequency distributions, and the significance of low-frequency and high-frequency components for time series modeling varies across domains(Zhang et al., 2024). As a result, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity. Existing pre-training frameworks (Dong et al., 2024; Nie et al., 2022; Lee et al., 2023), such as masked modeling and contrastive learning, were proposed to learn a unified representation from time domain. However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
@@
Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge. Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024). Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task (Chen et al., 2023a). However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) focus mainly on learning generalized representation during pre-training and overlook domain-specific representation. Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks. Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer. Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training. 2) adaptive use of domain-specific information in various downstream tasks.
@@
To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE). First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic information. We decompose individual time series using the Fourier transform with a novel frequency-based masking method, and then convert it back to the time domain to obtain decoupled time series for reconstruction. It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations. Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data. By setting up a register, we generate register tokens to learn each domain-specific information during pre-training. In a downstream scenario, the model adaptively selects Top-K vectors from the register that are close to the target domain of interest. During fine-tuning, we adjust the selected register tokens with a novel learnable low-rank matrix, which complements target-specific information to perform more flexible adaptive transfer. As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks. The contributions are summarized as follows:
  • We propose ROSE, a novel light weight general time series forecasting model using multi-domain datasets for pre-training and improving downstream fine-tuning performance and efficiency.
  • We propose a novel Decomposed Frequency Learning that employs multi-frequency masking to learn complex general temporal patterns from multi-domain data, empowering the model’s generalization capability.
  • We propose a novel TS-Register to capture domain-specific information in pre-training and enable adaptive transfer of target-oriented specific information for downstream tasks.
  • Our experiments with 7 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
@@ 2 Related work
2.1 Traditional time series forecasting
The statistical time series forecasting models like ARIMA (Box and Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity. With the rise of deep learning, many RNN-based models (Cirstea et al., 2019; Wen et al., 2017; Salinas et al., 2020) have been proposed, modeling the sequential data with an autoregressive process. CNN-based models (Luo and Wang, 2024; Liu et al., 2022a) have also received widespread attention due to their ability to capture local features. MICN (Wang et al., 2022) utilizes TCN to capture both local and global features, while TimesNet (Wu et al., 2022) focuses on modeling 2D temporal variations. However, both RNNs and CNNs struggle to capture long-term dependencies. Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction. However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
2.2 Pre-trained models for times seires forecasting
LLMs for time series forecasting: Recent studies have shown that Large Language Models (LLMs) can enhance time series forecasting with limited fine-tuning (Zhou et al., 2024), prompting (Jin et al., 2024; Cao et al., 2024), and modality alignment (Jin et al., 2024). GPT4TS (Zhou et al., 2024) fine-tunes a subset of LLM parameters, and achieves competitive performance by leveraging pre-trained text knowledge. TimeLLM (Jin et al., 2024) transforms time series into text to align with LLM representations. Despite their positive impact, LLMs’ high inference costs limit their practicality.
Time series foundation model: Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza and Mergenthaler-Canseco, 2023; Kamarthi and Prakash, 2023). MOMENT (Goswami et al., 2024) and MOIRAI (Woo et al., 2024) adopt a BERT-style pre-training approach, while Timer (Liu et al., 2024), Chronos (Ansari et al., 2024) and TimsFM (Das et al., 2023a) use a GPT-style pre-training approach, giving rise to improved performance in time series prediction. However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models. Different from previous approaches, ROSE pre-trains on large-scale data from various domains and it considers both generalized representations and domain-specific information, which facilitates flexible adaptive transfer in downstream tasks.
@@ 3.1 Architecture
The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning. First, ROSE is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction. We set up the reconstruction task to help the model understand time series comprehensively and set the prediction task to enhance the model’s few-shot and zero-shot abilities. Then, ROSE is fine-tuned with a target dataset in the downstream scenario.
As shown in Figure 2, ROSE employs the encoder-decoder architecture to model time series. The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies (Vaswani et al., 2017). The reconstruction decoder and prediction decoder use the same structure as the Transformer encoder. They are used for reconstruction and prediction tasks respectively. ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting (Nie et al., 2022).
Input Representations. To enhance the generalization of ROSE for adaptive transferring from multi-domains to different target domains, we model the inputs $\mathbf{x}$ with patch tokens and register tokens. Patch tokens are obtained by partitioning the time series using patching layers (Nie et al., 2022), to preserve local temporal information. Register tokens that capture domain-specific information will be introduced in Section 3.3.
@@ 3.2 Decomposed frequency learning
As shown in Figure 1, time series data are composed of multiple superimposed frequency components, resulting in the overlap of different temporal changes. Furthermore, low-frequency components typically contain information about overall trends and longer-scale variations, and high-frequency components usually contain information about short-term fluctuations and shorter-scale variations, therefore, understanding time series from low and high frequencies separately benefits general time series representation learning. Based on the observations above, we propose a novel frequency-based masked modeling that randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations. Finally, reconstruction task assists the model in comprehending the data from various frequency perspectives, enabling it to learn generalized representations. In contrast, existing frequency masking methods (Zhang et al., 2022), which randomly mask frequencies of a single time series once, show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
Figure 3: An illustration of decomposed frequency learning. Based on the sampled thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
Multi-frequency masking. As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham and Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
@@
To separately model high-frequency and low-frequency information in time series, we sample $K_{\mathrm{f}}$ thresholds ${\tau_{1},\tau_{2},\tau_{3},...,\tau_{K_{\mathrm{f}}}}$ and $K_{\mathrm{f}}$ random numbers ${\mu_{1},\mu_{2},\mu_{3},...,\mu_{K_{\mathrm{f}}}}$ for multi-frequency masks, where $\tau\in\mathrm{Uniform}(0,a)$, $a<L/2+1$ , and $\mu\in\mathrm{Bernoulli}(p)$. Each pair of $\tau_{i}$ and $\mu_{i}$ corresponds to the $i_{th}$ frequency mask. This generates a mask matrix $\mathbf{M}\in\{0,1\}^{K_{\mathrm{f}}\times(L/2+1)}$, where each row corresponds to the $i_{th}$ frequency mask, each column corresponds to the $j_{th}$ frequency, and each element $m_{ij}$ is 0 or 1, meaning that the $j_{th}$ frequency is masked with the $i_{th}$ frequency mask or not.
$m_{ij}=\left\{\begin{aligned} &\quad\mu_{i}&,if\,\ j<\tau_{i}\\ &(1-\mu_{i})&,if\,\ j>\tau_{i}\end{aligned}\right.,$ (2)
where $\tau_{i}$ and $\mu_{i}$ denote the threshold and random number for the $i_{th}$ frequency domain mask. If $\mu_{i}=1$, it means that frequency components above $\tau_{i}$ will be masked, indicating to mask high frequency , as shown by the threshold $\tau_{1}$ in Figure 3. Conversely, if $\mu_{i}=0$, it signifies that frequency components below $\tau_{i}$ will be masked, indicating to mask low frequency, exemplified by threshold $\tau_{2}$ in Figure 3.
After obtaining the mask matrix $\mathbf{M}$, we replicate $\mathbf{x}_{\mathrm{freq}}$ $K_{\mathrm{f}}$ times to get the $\mathbf{X}_{\mathrm{freq}}\in\mathbb{C}^{K_{\mathrm{f}}\times L/2+1}$ and perform element-wise Hadamard product with the mask matrix $\mathbf{M}$ to get masked frequency of time series. Then, we use the inverse Real Fast Fourier Transform (irFFT) to convert the results from the frequency domain back to the time domain and get $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}=\{\mathbf{x}_{\mathrm{mask}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$, where each $\mathbf{x}_{\mathrm{mask}}^{i}\in\mathbb{R}^{L}$ corresponding to masking with a different threshold $\tau_{i}$.
@@
Representation learning. As shown in the yellow part of Figure 3, after obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token. We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3. Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the Transformer encoder to obtain the representation of each masked series. These representations are then aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$. The aggregator is the averaging operation.
$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
Reconstruction task. After obtaining the representation $\mathbf{S}_{\mathrm{m}}$, we feed it into the reconstruction decoder, which shares same stucture as the Tranformer encoder, and ultimately reconstruct the original sequence $\mathbf{\hat{x}}\in\mathbb{R}^{L}$ through the reconstruction head, which is a linear layer. As frequency domain masking affects the overall time series, we compute the Mean Squared Error (MSE) reconstruction loss for the entire time series.